The width of a swimming pool is 12 feet less than the pool's length. The area of the pool is 540 square feet. What are the dimensions of the pool?

Found 3 solutions by Boreal, ikleyn, MaxWong:
Answer by Boreal(15235)

width=x
length=x+12, the width is 12 feet less, the length is 12 feet more. I could make length x and width x-12, but I'd rather deal with + than -. You get the same answer.
x(x+12)=540. LxW=area
x^2+12x-540=0
(x+30)(x-18)=0
x=-30, 18. Only the second answer makes sense
width is 18 feet and length is 30 feet.

Answer by ikleyn(52781)

.
30 ft long and 18 ft wide.


Answer by MaxWong(38)

x(x-12)=540
x^2 - 12x - 540 = 0
(x-30)(x+18)=0
x = 30 or x = -18(rejected)
Length = 30 and Width = 30 - 12 = 18